# Diverse Part Synthesis for 3D Shape Creation

### v1

@@ 1 Introduction
Synthesizing 3D shapes is an important research problem in computer graphics, which has applications in the creation of content for games, simulations, and virtual worlds. The latest methods for automatic or semi-automatic synthesis of 3D shapes make use of deep neural networks, generating shapes either through the sampling of latent or noise vectors chen2019learning ; li2021sp , diffusion models zeng2022lion ; hui2022neural , or through language guidance chen2018text2shape ; liu2022towards ; sanghi2022clip . Many of these methods synthesize shapes as a whole in the form of point clouds li2021sp ; zeng2022lion , volumes sanghi2022clip , or implicit functions chen2019learning ; hui2022neural ; li2022learning .
On the other hand, shapes can also be represented as a composition of parts, either in the form of a graph or hierarchy that explicitly encodes the shape’s structure mitra2013structure . This representation has advantages in modeling and analysis, such as the possibility of modifying and exchanging individual parts or annotating shapes with semantic information. A recent line of work has used part-based representations for 3D shape synthesis with neural networks, proposing methods that can be employed for a variety of useful tasks such as shape synthesis from latent representations schor2019componet ; mo2019structurenet ; wu2019sagnet ; wu2020pq ; niu2022rim , part-based shape reconstruction dubrovina2019composite ; petrov2023anise , and shape interpolation yang2022dsg ; wu2020pq ; mo2019structurenet .
@@
Despite the various uses of these methods, one limitation of the current work is that it is difficult to control the generation of individual shape parts. That is, if the user would like to modify a specific part that has been generated, the existing methods do not provide a specific mechanism for generating multiple meaningful suggestions to replace the part. Such a functionality is ideal in an iterative modeling framework where the user starts with an initial shape suggested by the modeling system and refines its details to progressively approximate a mental design. In the current methods, shapes are either represented as latent vectors that encode entire shapes wu2020pq ; niu2022rim or do not offer mechanisms for regenerating suitable latent codes for individual parts while ensuring the variety of the parts and their consistency with the current shape schor2019componet ; dubrovina2019composite ; petrov2023anise . A few works have introduced methods for disentangling parts from structure yang2022dsg , which allow interpolating geometry or structure independently. However, control and resampling of individual parts is not explicitly enabled by the disentangled representation.
In this paper, we investigate the development of a method that synthesizes shapes in a part-based representation while allowing the user to generate a variety of distinct suggestions for individual parts. Specifically, we perform a comparative study of multimodal deep generative models that enable to synthesize multiple, diverse part suggestions to compose a shape. The use of a multimodal generative model is especially important to ensure the diversity of the suggested parts. Deep generative models are known for facing problems such as mode collapse, which results in the networks operating practically in a deterministic manner, where highly similar outputs are generated for a same input. Therefore, in this work, we experiment with various multimodal generative models and conduct a comparative study of these models, identifying the most proficient ones in generating suggested parts with both shape fidelity and diversity. Specifically, we evaluate mixture density networks (MDNs) bishop1994mixture , as well as conditional generative models implemented with generative adversarial networks (GANs) goodfellow2014generative , generator networks trained with implicit maximum likelihood estimation (IMLE) li2018implicit , and denoising diffusion probabilistic models (DDPMs) ho2020denoising .
@@
The incremental shape synthesis is enabled by a method composed of three main neural networks: (i) a decoder that generates shapes in an implicit representation based on a latent vector, (ii) a spatial transformer jaderberg2015spatial that positions the synthesized parts in their target locations, and (iii) a multimodal generative model that suggests latent vectors of suitable parts that can be connected to a given partial assembly. In addition, an autoencoder (AE) architecture ensures that a latent space of shape parts is learned by the method.
@@
  • We perform a comparative study of multimodal generative models for suggesting diverse parts for shape synthesis and demonstrate that the conditional models implemented with IMLE and DDPMs provide the best overall part suggestions in terms of visual quality and diversity;
@@ 2.1 Part-based shape synthesis
Shape synthesis by data-driven part assembly has been an area of active research in computer graphics for a few decades, given the flexibility of modeling shapes as a composition of parts funkhouser2004modeling ; mitra2013structure . Earlier part-based assembly methods make use of a variety of techniques such as part suggestion via probabilistic reasoning chaudhuri2011probabilistic ; kalogerakis2012probabilistic ; sung2017complementme or set evolution xu2012fit ; guan2022fame . These methods typically start from a predefined set of parts and combine the parts to generate new shapes. Recent methods for part-based synthesis make use of deep neural networks, encoding shapes and their parts in a variety of representations, such as point clouds schor2019componet , voxel sets dubrovina2019composite ; li2020learning , and implicit fields wu2020pq ; petrov2023anise . Part-based assembly methods based on neural networks can be roughly classified into four groups.
@@
The first group of methods assume that a set of parts is given as input and mainly provide recombinations of these parts for generating new shapes, similarly as the earlier part-based assembly methods. For example, in their ComplementMe system, Sung et al. sung2017complementme pose the problem of suggesting parts to be added to an incomplete partial assembly as a retrieval problem solved with a neural network. Similarly, Sung et al. sung2018learning encode the interchangeability or complementarity of shape parts in a dual embedding, while Huang et al. huang2020generative introduce a method for reassembling parts through dynamic graph learning. Although these methods can be used for generating shapes with new combinations of parts, the geometry of the parts is limited to the existing collection, and thus parts with varied geometry cannot be generated.
@@
The second group of methods synthesizes the part structure and geometry of the parts independently, thus enabling the generation of novel shape structures and their parts. This is typically achieved with the combination of a part generation network and a part placement network, which also inspired the architecture of our method. For example, Dubrovina et al. dubrovina2019composite introduce a neural network for part-based reconstruction based on a part decoder and localization network. Schor et al. schor2019componet propose a method composed of a part decoder network and a composition network, generating shapes in a part-based representation by randomly sampling part and composition latent codes, while Li et al. li2020learning also use latent codes for generating parts as volumes but predict part placement from the synthesized volumes. ANISE petrov2023anise reconstructs shapes with a part-based reconstruction pipeline based on part and position codes and an implicit representation of parts. Although the methods in this group allow generating or reconstructing shapes in a part-based presentation, they provide limited or no functionality for easily replacing or regenerating specific shape parts. All parts are either assumed to be independent and fully exchangeable schor2019componet , or the methods do not provide mechanisms for regenerating specific parts codes from partial shapes, given their main goal of shape reconstruction dubrovina2019composite ; petrov2023anise .
@@
Moreover, the third group of methods synthesize the part structure and individual parts of a shape based on a single latent code. The method either generates the entire shape from the vector or iteratively generates a sequence of vectors from the initial one to guide the generation of parts. This is achieved with a hierarchical graph representation in StructureNet mo2019structurenet , a joint latent space of structure and geometry in SAGNet wu2019sagnet , sequential generation of a part structure in PQ-NET wu2020pq , or a sequence of latent spaces encoded from a shape’s hierarchy in LSD-StructureNet roberts2021lsd . These methods are quite versatile given that a complex shape can be generated from a single latent code. However, since individual parts are not associated to specific entries of the latent codes, the regeneration of specific parts dependent on other parts of the shape cannot be easily achieved.
@@
Finally, the fourth group of methods propose to learn part-based representations in an unsupervised or self-supervised manner. RIM-Net niu2022rim learns recursive implicit fields for encoding shapes in a hierarchical structure. DSG-Net yang2022dsg encodes shapes in a representation that disentangles geometry and structure, such that the structure and geometry of shapes can be sampled or interpolated independently. SPAGHETTI hertz2022spaghetti supports direct editing of 3D models by learning a representation of shape parts and their transformations. The advantage of these methods is that labeled data is not required for training, since the loss ensures that a part decomposition is obtained in an unsupervised manner. However, the part segmentation discovered by the methods may not be ideal for certain shapes, and the part decomposition is not guaranteed to fit the user’s intended semantic abstraction of the shapes.
@@ 2.2 Deep generative models with multimodality
A common problem in machine learning is how to synthesize results with diversity, given that common generative models operate in practice in a monomodal manner, synthesizing nearly identical outputs for a given input. For example, research in image translation has reported this effect in conditional image generation isola2017image . In the case of GANs, the monomodal generation arises due to the problem of mode collapse, where the training process maps a certain group of inputs to a single output that should be predicted.
Research in this area has attempted to address the diversity problem in different manners. One group of methods predict a distribution of latent vectors, rather than a single vector, for a given input. For example, the distribution can be encoded in the form of a mixture of Gaussians bishop1994mixture , so that multiple, distinct vectors can be sampled from the distribution. This type of approach has been used for shape generation sung2017complementme ; jahan2021semantics . Since the distribution is learned independently from the embedding, it is not guaranteed to be contained in the latent manifold, which implies that decoding samples of the distribution may lead to suboptimal results.
Another line of work introduces conditional generative models for multimodal generation, e.g., conditioning the input to a GAN mirza2014conditional or to a variational autoencoder (VAE) sohn2015learning . Wu et al. wu2020multimodal combine a VAE with a GAN for multimodal generation in shape completion. However, despite the multimodal output, the quality of the results suffers from the VAE encoding and decoding. To prevent mode collapse incurred by GANs, IMLE has been proposed by Li and Malik li2018implicit . The main idea of this method is to match each training sample to a generated sample during training, so that the network takes into consideration all possible modes in the training set. This helps to alleviate the problem of mode collapse when learning generative models. Recently, DDPMs proposed by Ho et al. ho2020denoising have emerged as a novel powerful deep generative model. DDPMs can be made conditional by using the condition information as an additional input through the diffusion pathway nichol2021improved .
The sampling procedure of the aforementioned conditional models is based the idea of an implicit generative model mohamed2016learning , i.e., adding a random noise vector $\mathbf{n}\sim\mathcal{N}(0,\mathbf{I})$ to a deterministic parameterized transformation $F_{\theta}$, so that, given the conditional input $\mathbf{c}$, the predicted sample $\hat{\mathbf{x}}$ is drawn from $F_{\theta}$ as:
@@
Conditional models have been applied to tasks such as super-resolution chen2018fsrnet ; li2020multimodal ; rombach2022high ; saharia2023image and image synthesis from a given layout isola2017image ; li2019diverse ; li2020multimodal ; rombach2022high ; wang2022semantic and also to shape completion wu2020multimodal ; zhou20213d ; arora2022multimodal ; cheng2023sdfusion ; li2023diffusion ; chu2023diffcomplete and shape generation from text chen2018text2shape ; liu2022towards ; zeng2022lion ; cheng2023sdfusion ; li2023diffusion . Note that these approaches predict or complete entire shapes at a time.
@@ 3 Shape synthesis with part suggestions
In this section, we introduce our method for synthesizing 3D shapes in a part-based representation, which we use for evaluating the multimodal part suggestion techniques. Our method is composed of two main deep neural networks: (i) a part composition network (PCN) that learns a latent space of shape parts and the affine transformation applied to each part and (ii) a part suggestion network (PSN) that learns to generate the latent codes of the possible parts given an incomplete partial assembly. The PSN suggests parts with one of the multimodal generative models that we cover in our study. Moreover, to achieve a higher visual quality of the synthesized shapes, we use an implicit decoder chen2019learning , which decodes latent vectors generated by the PSN into implicit fields representing shape parts. We illustrate in Figure 2 the architecture of the PCN and PSN and the process of part suggestion and placement.
Given an initial shape part or partial assembly as input, the PSN generates possible latent codes representing the shape parts that can complement the current partial assembly. Then, given a part suggestion, we transform the decoded shape part into its target position based on the prediction of affine transformations learned by the PCN. Thus, to generate a complete shape, we start from an initial part and iteratively add more parts to the assembly based on the suggestions from the PSN and decoding and placement of the PCN. This process can be carried out in an interactive manner, where a user selects one of the suggested parts at each iteration, or in a randomized manner, where one or multiple suggestions are automatically selected to compose a set of randomly sampled shapes.

### v2

@@ 1 Introduction
Synthesizing 3D shapes is an important research problem in computer graphics, which has applications in the creation of content for games, simulations, and virtual worlds. The latest methods for automatic or semi-automatic synthesis of 3D shapes make use of deep neural networks, generating shapes either through the sampling of latent or noise vectors [1, 2], diffusion models [3, 4], or through language guidance [5, 6, 7]. Many of these methods synthesize shapes as a whole in the form of point clouds [2, 3], volumes [7], or implicit functions [1, 4, 8].
On the other hand, shapes can also be represented as a composition of parts, either in the form of a graph or hierarchy that explicitly encodes the shape’s structure [9]. This representation has advantages in modeling and analysis, such as the possibility of modifying and exchanging individual parts or annotating shapes with semantic information. A recent line of work has used part-based representations for 3D shape synthesis with neural networks, proposing methods that can be employed for a variety of useful tasks such as shape synthesis from latent representations [10, 11, 12, 13, 14], part-based shape reconstruction [15, 16], and shape interpolation [17, 13, 11].
@@
Despite the various uses of these methods, one limitation of the current work is that it is difficult to control the generation of individual shape parts. That is, if the user would like to modify a specific part that has been generated, the existing methods do not provide a specific mechanism for generating multiple meaningful suggestions to replace the part. Such a functionality is ideal in an iterative modeling framework where the user starts with an initial shape suggested by the modeling system and refines its details to progressively approximate a mental design. In the current methods, shapes are either represented as latent vectors that encode entire shapes [13, 14] or do not offer mechanisms for regenerating suitable latent codes for individual parts while ensuring the variety of the parts and their consistency with the current shape [10, 15, 16]. A few works have introduced methods for disentangling parts from structure [17], which allow interpolating geometry or structure independently. However, control and resampling of individual parts is not explicitly enabled by the disentangled representation.
In this paper, we investigate the development of a method that synthesizes shapes in a part-based representation while allowing the user to generate a variety of distinct suggestions for individual parts. Specifically, we perform a comparative study of multimodal deep generative models that enable to synthesize multiple, diverse part suggestions to compose a shape. The use of a multimodal generative model is especially important to ensure the diversity of the suggested parts. Deep generative models are known for facing problems such as mode collapse, which results in the networks operating practically in a deterministic manner, where highly similar outputs are generated for a same input. Therefore, in this work, we experiment with various multimodal generative models and conduct a comparative study of these models, identifying the most proficient ones in generating suggested parts with both shape fidelity and diversity. Specifically, we evaluate mixture density networks (MDNs) [18], as well as conditional generative models implemented with generative adversarial networks (GANs) [19], generator networks trained with implicit maximum likelihood estimation (IMLE) [20], and denoising diffusion probabilistic models (DDPMs) [21].
@@
The incremental shape synthesis is enabled by a method composed of three main neural networks: (i) a decoder that generates shapes in an implicit representation based on a latent vector, (ii) a spatial transformer [22] that positions the synthesized parts in their target locations, and (iii) a multimodal generative model that suggests latent vectors of suitable parts that can be connected to a given partial assembly. In addition, an autoencoder (AE) architecture ensures that a latent space of shape parts is learned by the method.
@@
  • We perform a comparative study of multimodal generative models for suggesting diverse parts for shape synthesis;
  • We demonstrate that the conditional models implemented with IMLE and DDPMs provide the best overall part suggestions in terms of visual quality and diversity;
@@ 2.1 Part-based shape synthesis
Shape synthesis by data-driven part assembly has been an area of active research in computer graphics for a few decades, given the flexibility of modeling shapes as a composition of parts [23, 9]. Earlier part-based assembly methods make use of a variety of techniques such as part suggestion via probabilistic reasoning [24, 25, 26] or set evolution [27, 28]. These methods typically start from a predefined set of parts and combine the parts to generate new shapes. Recent methods for part-based synthesis make use of deep neural networks, encoding shapes and their parts in a variety of representations, such as point clouds [10], voxel sets [15, 29], and implicit fields [13, 16]. Part-based assembly methods based on neural networks can be roughly classified into four groups.
@@
The first group of methods assume that a set of parts is given as input and mainly provide recombinations of these parts for generating new shapes, similarly as the earlier part-based assembly methods. For example, in their ComplementMe system, Sung et al. [26] pose the problem of suggesting parts to be added to an incomplete partial assembly as a retrieval problem solved with a neural network. Similarly, Sung et al. [30] encode the interchangeability or complementarity of shape parts in a dual embedding, while Huang et al. [31] introduce a method for reassembling parts through dynamic graph learning. To improve the visual fidelity of the assembled parts, Yin et al. [32] reconstruct the part joints as implicit surfaces. Although these methods can be used for generating shapes with new combinations of parts, the geometry of the parts is limited to the existing collection, and thus parts with varied geometry cannot be generated.
@@
The second group of methods synthesizes the part structure and geometry of the parts independently, thus enabling the generation of novel shape structures and their parts. This is typically achieved with the combination of a part generation network and a part placement network, which also inspired the architecture of our method. For example, Dubrovina et al. [15] introduce a neural network for part-based reconstruction based on a part decoder and localization network. Schor et al. [10] propose a method composed of a part decoder network and a composition network, generating shapes in a part-based representation by randomly sampling part and composition latent codes, while Li et al. [29] also use latent codes for generating parts as volumes but predict part placement from the synthesized volumes. ANISE [16] reconstructs shapes with a part-based reconstruction pipeline based on part and position codes and an implicit representation of parts. Although the methods in this group allow generating or reconstructing shapes in a part-based presentation, they provide limited or no functionality for easily replacing or regenerating specific shape parts. All parts are either assumed to be independent and fully exchangeable [10], or the methods do not provide mechanisms for regenerating specific parts codes from partial shapes, given their main goal of shape reconstruction [15, 16].
@@
Moreover, the third group of methods synthesize the part structure and individual parts of a shape based on a single latent code. The method either generates the entire shape from the vector or iteratively generates a sequence of vectors from the initial one to guide the generation of parts. This is achieved with a global-to-local GAN alongside a part refiner by Wang et al. [33], a hierarchical graph representation in StructureNet [11], a joint latent space of structure and geometry in SAGNet [12], sequential generation of a part structure in PQ-NET [13], or a sequence of latent spaces encoded from a shape’s hierarchy in LSD-StructureNet [34]. These methods are quite versatile given that a complex shape can be generated from a single latent code. However, since individual parts are not associated to specific entries of the latent codes, the regeneration of specific parts dependent on other parts of the shape cannot be easily achieved.
@@
Finally, the fourth group of methods propose to learn part-based representations in an unsupervised or self-supervised manner. RIM-Net [14] learns recursive implicit fields for encoding shapes in a hierarchical structure. DSG-Net [17] encodes shapes in a representation that disentangles geometry and structure, such that the structure and geometry of shapes can be sampled or interpolated independently. SPAGHETTI [35] supports direct editing of 3D models by learning a representation of shape parts and their transformations. The advantage of these methods is that labeled data is not required for training, since the loss ensures that a part decomposition is obtained in an unsupervised manner. However, the part segmentation discovered by the methods may not be ideal for certain shapes, and the part decomposition is not guaranteed to fit the user’s intended semantic abstraction of the shapes.
@@ 2.2 Deep generative models with multimodality
A common problem in machine learning is how to synthesize results with diversity, given that common generative models operate in practice in a monomodal manner, synthesizing nearly identical outputs for a given input. For example, research in image translation has reported this effect in conditional image generation [36]. In the case of GANs, the monomodal generation arises due to the problem of mode collapse, where the training process maps a certain group of inputs to a single output that should be predicted.
Research in this area has attempted to address the diversity problem in different manners. One group of methods predict a distribution of latent vectors, rather than a single vector, for a given input. For example, the distribution can be encoded in the form of a mixture of Gaussians [18], so that multiple, distinct vectors can be sampled from the distribution. This type of approach has been used for shape generation [26, 37]. Since the distribution is learned independently from the embedding, it is not guaranteed to be contained in the latent manifold, which implies that decoding samples of the distribution may lead to suboptimal results.
Another line of work introduces conditional generative models for multimodal generation, e.g., conditioning the input to a GAN [38] or to a variational autoencoder (VAE) [39]. Wu et al. [40] combine a VAE with a GAN for multimodal generation in shape completion. However, despite the multimodal output, the quality of the results suffers from the VAE encoding and decoding. To prevent mode collapse incurred by GANs, IMLE has been proposed by Li and Malik [20]. The main idea of this method is to match each training sample to a generated sample during training, so that the network takes into consideration all possible modes in the training set. This helps to alleviate the problem of mode collapse when learning generative models. Recently, DDPMs proposed by Ho et al. [21] have emerged as a novel powerful deep generative model. DDPMs can be made conditional by using the condition information as an additional input through the diffusion pathway [41].
The sampling procedure of the aforementioned conditional models is based the idea of an implicit generative model [42], i.e., adding a random noise vector $\mathbf{n}\sim\mathcal{N}(0,\mathbf{I})$ to a deterministic parameterized transformation $F_{\theta}$, so that, given the conditional input $\mathbf{c}$, the predicted sample $\hat{\mathbf{x}}$ is drawn from $F_{\theta}$ as:
@@
Conditional models have been applied to tasks such as super-resolution [43, 44, 45, 46] and image synthesis from a given layout [36, 47, 44, 45, 48] and also to shape completion [49, 40, 50, 51, 52, 53, 54] and shape generation from text [5, 6, 3, 52, 53]. Note that these approaches predict or complete entire shapes at a time.
@@ 3 Shape synthesis with part suggestions
In this section, we introduce our method for synthesizing 3D shapes in a part-based representation, which we use for evaluating the multimodal part suggestion techniques. Our method is composed of two main deep neural networks: (i) a part composition network (PCN) that learns a latent space of shape parts and the affine transformation applied to each part and (ii) a part suggestion network (PSN) that learns to generate the latent codes of the possible parts given an incomplete partial assembly. The PSN suggests parts with one of the multimodal generative models that we cover in our study. Moreover, to achieve a higher visual quality of the synthesized shapes, we use an implicit decoder [1], which decodes latent vectors generated by the PSN into implicit fields representing shape parts. We illustrate in Figure 2 the architecture of the PCN and PSN and the process of part suggestion and placement.
We first train the PCN to derive the latent space of shape parts and then use the latent codes of shape parts to train the PSN for generating multimodal part suggestions. The details of training the PCN and PSN are discussed in the following subsections. Once all the network modules of the PCN and PSN are trained, new shapes can be synthesized iteratively with incrementally suggested shape parts. Given an initial shape part or partial assembly as input, the PSN generates possible latent codes representing the shape parts that can complement the current partial assembly. Then, given a part suggestion, we transform the decoded shape part into its target position based on the prediction of affine transformations learned by the PCN. Thus, to generate a complete shape, we start from an initial part and iteratively add more parts to the assembly based on the suggestions from the PSN and decoding and placement of the PCN. This process can be carried out in a user-in-the-loop manner, where a user selects one or more of the suggested parts at each iteration, or in a randomized manner, where one or multiple suggestions are automatically selected to compose a set of randomly sampled shapes. We detail this iterative shape synthesis process via incremental part suggestion in Algorithm 1. The synthesis is carried out by the function IterativeSynthesis($\mathbf{P}$,$K$), which takes an initial shape part $\mathbf{P}$ and a predefined number of iterations $K$ as input and outputs a set $\mathcal{A}$ containing the final synthesized shapes (assemblies). In Algorithm 1, each of the functions, i.e., Encoder, Decoder, LocalizationNetwork, Transform, and ImplicitDecoder, corresponds to a network module presented in Figure 2 (in gray), and UserSelection represents the process of users’ visual inspection and selection of one or multiple plausible partial assemblies at the end of each iteration.
Input: $\mathbf{P}$, $K$
Output: $\mathcal{A}$
1 function IterativeSynthesis($\mathbf{P}$,$K$)
2       $\mathcal{A}\leftarrow\{\mathbf{P}\}$
3       $i\leftarrow 1$
4       while $i\leq K$ do
5             foreach $\mathbf{A}\in\mathcal{A}$ do
6                   $\mathcal{A}^{\prime}\leftarrow\emptyset$
7                   $\mathbf{y}\leftarrow\texttt{Encoder($\mathbf{A}$)}$
8                   $\hat{\mathcal{Z}}\leftarrow\texttt{MultimodalGenerativeModel($\mathbf{y}$)}$
9                   foreach $\hat{\mathbf{z}}\in\hat{\mathcal{Z}}$ do
10                         $\hat{\mathbf{P}}\leftarrow\texttt{Decoder($\hat{\mathbf{z}}$)}$
11                         $\hat{s},\hat{\mathbf{t}}\leftarrow\texttt{LocalizationNetwork($\hat{\mathbf{P}}$)}$
12                         $\hat{\mathbf{P}}\leftarrow\texttt{ImplicitDecoder($\hat{\mathbf{z}}$)}$
13                         $\hat{\mathbf{P}}^{\prime}\leftarrow\texttt{Transform($\hat{\mathbf{P}}$,$\hat{s}$,$\hat{\mathbf{t}}$)}$
14                         $\mathbf{A}^{\prime}\leftarrow\mathbf{A}+\hat{\mathbf{P}}^{\prime}$
15                         $\mathcal{A}^{\prime}\leftarrow\mathcal{A}^{\prime}\cup\{\mathbf{A}^{\prime}\}$
16
17                   end foreach
18
19             end foreach
20            $\mathcal{A}\leftarrow\texttt{UserSelection($\mathcal{A}^{\prime}$)}$
21             $i\leftarrow i+1$
22
23       end while
24      return $\mathcal{A}$
25
26 end
Algorithm 1 Iterative synthesis